Can someone give a good explanation?

A spherical wave travels outward from a point source What is the phase difference between the two points on the wave marked with dots? A. B. C. D. E. π/4 radians π /2 radians π radians 7 T /2 radians 7 π radians
Answer #1

The correct answer is C.

In the given problem, a point source is giving rise to wavefronts that are spherical in shape. The wavefronts are emitted in all directions and the energy from the source is transmitted through them. Wavefronts are the surfaces over which the oscillations due to waves are the same. This also means that the phase difference between two points on the same wavefront is zero. In the case of particles oscillating in two different and consecutive wavefronts, the phase difference between them is 2\pi. Moreover, the distance between two consecutive wavefronts is called the wavelength of the emitted wave.

In the given picture, one of the points is on the wavefront while other point is in between the two wavefronts. In this case, the two points are almost half wavelength apart. Since phase difference between oscillating particles which are a wavelength apart is 2\pi, the phase difference for particles half wavelength apart is \pi.